# VertiBench: Advancing Feature Distribution Diversity in Vertical Federated Learning Benchmarks

### v1

@@ 4.2 Feature Importance
Consider a feature index set $\mathcal{A}=\{1,2,...,m\}$ and a characteristic function $v:2^{\mathcal{A}}\rightarrow\mathbb{R}$ such that $v(\emptyset)=0$. Let $\phi_{j}(v)$ denote the importance of the $j$-th feature on $v$ such that $\sum_{j=1}^{m}\phi(j)=v(\mathcal{A})$. Assume that the indices in $\mathcal{A}$ are randomly distributed to $K$ parties with probabilities $r_{1},...,r_{K}$ where $\sum_{i=1}^{K}{r_{i}}=1$. Given budgets $b_{i}=r_{i}v(\mathcal{A})$, let $Z_{i}$ be the sum of feature importance for party $i$. Then, we have $\forall i\in[1,K],\mathbf{E}[Z_{i}]=b_{i}$ and $\mathbf{E}[Z_{i}]\propto r_{i}$.
@@
The proof of Theorem 1 is provided in Appendix A. The metric of importance, $\phi_{j}(v)$, comprises the Shapley value and the recently proposed Shapley-CMI Han et al. (2021). We assert in Theorem 1 that the expected cumulative importance of each party is proportional to the ratio generated by the Dirichlet distribution. The inherent properties of the Dirichlet distribution ensure that: (1) a larger value of $\alpha_{i}$ leads to a higher expected value of $r_{i}$, and (2) a smaller value of $\|\{\alpha_{i}\}_{i=1}^{K}\|_{2}$ results in a larger variance in $r_{i}$. Hence, the proposed method naturally aligns with the requirements for feature importance.
4.3 Feature Correlation
In the initial stages of our investigation into feature-split methods based on correlation, we first look at the evaluation of feature correlation. Building upon established methods that utilize a metric grounded in correlation matrices Taylor (2020); Wang and Zheng (2014), we propose a novel metric to examine the correlation when the parties involved possess unequal numbers of features. Our approach hinges on the use of the standard variance of the singular values of the correlation matrix. This serves as an efficient measure of the overall correlation between two parties. Since the feature-wise correlation is an orthogonal research area, we selected Spearman rank correlation Zar (2005) due to its capability to handle non-linear correlation.
To elaborate further, we denote the column-wise correlation matrix between two matrices, $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$, as $\text{cor}(\mathbf{X}_{i},\mathbf{X}_{j})$. As a result, we formally define the correlation between two entities, $\mathbf{X}_{i}\in\mathbb{R}^{n\times m_{i}}$ and $\mathbf{X}_{j}\in\mathbb{R}^{n\times m_{j}}$, in terms of their respective parties as Equation 2.
$\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j}):=\frac{1}{d}\sqrt{\sum_{i=1}^{d}{\left(\sigma_{i}(\text{cor}(\mathbf{X}_{i},\mathbf{X}_{j}))-\overline{\sigma}\right)^{2}}},\quad d=\min(m_{i},m_{j})$ (2)
In this equation, $\sigma_{i}(\cdot)$ means the $i$-th singular value of a matrix, while $\overline{\sigma}$ stands for their mean value.
$\text{Icor}(\mathbf{X}_{1},\dots,\mathbf{X}_{K}):=\frac{1}{K(K-1)}\sum_{i=1}^{K}\sum_{j=1,j\neq i}^{K}{\text{Pcor}(\mathbf{X}_{i},\mathbf{X}_{j})}$ (3)
@@ 4.3 Feature Correlation
This correlation-based feature-split algorithm, as depicted in Algorithm 1, is meticulously designed to allocate features across multiple parties while taking into account the correlations inherent among the features. The algorithm’s operation is premised on the provision of a defined number of features for each party, represented as $m_{1},\dots,m_{K}$. Commencing with the initialization of a column permutation matrix, denoted as $\mathbf{P}$, to an identity matrix (line 1), the algorithm proceeds to define a score function, $f(\mathbf{P};\mathbf{X})$, which represents the overall correlation Icor after the features have undergone permutation by $\mathbf{P}$ (line 2). Subsequently, the algorithm determines the lower and upper bound of the score function (lines 3-4). This forms the basis for calculating the target correlation $f^{*}(\mathbf{X};\beta)$, which is a linear interpolation between the lower and upper bounds controlled by the correlation index $\beta$ (line 5). Next, the algorithm locates the optimal permutation matrix $\mathbf{P}^{*}$ by solving an permutation-based optimization problem. Notably, we employ the Biased Random-Key Genetic Algorithm (BRKGA) Gonçalves and Resende (2011) for this purpose. The final step of the algorithm splits the features according to the derived optimal permutation and the pre-set number of features for each party (lines 6-7).
@@
Owing to the fact that the optimization approach requires many invocations of Icor, it is important that this process is conducted with the highest degree of efficiency. For datasets of smaller dimensions, singular values can be directly computed utilizing Singular Value Decomposition (SVD) Baker (2005). However, in the case of high-dimensional datasets, we resort to employing Truncated SVD Hansen (1990) to estimate the largest top-$d_{t}$ singular values, with the remaining singular values assumed as zero prior to calculating the standard variance. It is worth noting that we make use of GPU acceleration to expedite the computation of Icor, thereby ensuring that the optimization procedure is as swift and efficient as possible. Our experiments, as presented in Appendix D, validate that both split methods can complete within a reasonable time.
Empirical Validation. We conduct extensive experiments to rigorously evaluate the practical performance of our proposed correlation evaluation metric and the correlation-based feature-split algorithm; the details are in Appendix D. Briefly, for the correlation evaluation metric Icor, we observe that Pcor mirrors the behavior of mcor Taylor (2020) in assessing inner-party correlation and displays a similar trend to mcor for inter-party correlation evaluation. Moreover, we split features of synthetic datasets of different $\beta$ values using Algorithm 1, contrasting it with a random split. The absolute correlation matrix visualized in Figure 1 suggests that as $\beta$ increases, so does inter-party correlation. In contrast, random feature splitting does not effectively portray scenarios with low inter-party correlation.
(a) corr-split: $\beta=0$
(b) corr-split: $\beta=0.5$
(c) corr-split: $\beta=1.0$
(d) random-split
Figure 1: Absolute correlation matrix of the global dataset with party boundaries indicated by red lines. Icor means inter-party correlation. (a),(b),(c) - correlation-based split; (d) random split.
5 Real-world VFL Datasets
Real-world VFL datasets, though highly desirable, are limited in scope and type, often encompassing tabular-tabular data, as in Vehicle Duarte and Hu (2004), Movielens Hu et al. (2022), Songs Hu et al. (2022), and tabular-image data, as in NUS-WIDE Chua et al. (July 8-10, 2009). Notably missing are image-image datasets. Addressing this, we introduce a real-world VFL dataset, Satellite, adapted from Cornebise et al. (2022), containing 62,832 images across 16 parties, simulating a practical VFL scenario of collaborative location identification via multiple satellites. Further details on Satellite’s construction are in Appendix E.
An in-depth analysis of the Satellite dataset, using our proposed metrics and the visualization of absolute correlation matrix (Figure 1(b)), reveals low inter-party correlation (Icor) and high inner-party correlation, similar to NUS-WIDE (Figure 1(a)). These results underscore the fact that random feature splits, which usually result in larger $\beta$ values (Figures 0(d) and 1(c)), may not truly represent real-world scenarios, reinforcing the need for systematic VFL datasets generation methods.
Estimating $\alpha$ and $\beta$ for real VFL datasets. In order to align real datasets with the synthetic ones generated by VertiBench, we put forward methods to estimate $\alpha$ and $\beta$ for real VFL datasets.
To calculate $\alpha$, we determine the significance of each party by adding up the Shapley value of its features. We do this efficiently by estimating Shapley values on a select subset. These Shapley values are then normalized and treated as Dirichlet parameters $\alpha_{i}$ for each party $P_{i}$, in line with Theorem 1. To approximate the scale of the Dirichlet parameters and align them with the generation of synthetic datasets, we find a symmetric Dirichlet distribution $\text{Dir}(\overline{\alpha})$ that has the same variance as $\text{Dir}(\alpha_{1},\dots,\alpha_{K})$, as given in Proposition 2. This value of $\overline{\alpha}$ reflects the variance in feature importance across parties. The proof is provided in Appendix A.
@@ 5 Real-world VFL Datasets
Using the estimated $\alpha$ and $\beta$, we display the importance and correlation of existing real datasets within the VertiBench-supported range in Figure 1(c). We note that real datasets represent a limited set of VFL scenarios with a large $\alpha$ and small $\beta$, indicating a high degree of feature imbalance and low inter-party correlation. Further, conducting random feature splits, as is common in existing VFL experiments, results in a distinct extreme characterized by high values of both $\alpha$ and $\beta$. This observation underscores the importance of VertiBench, which can generate a broad range of VFL scenarios for robust evaluation of VFL algorithms.
(a) NUS-WIDE correlation
(b) Satellite correlation
(c) Estimated $\alpha$ and $\beta$
Figure 2: (a), (b) Matrix of feature absolute correlation with party boundaries marked in red; (c) scope coverage of real and randomly split datasets in terms of feature importance and correlation.
@@ 6 Experiment
This section comprehensively benchmarks cutting-edge VFL algorithms. The experimental settings are delineated in Section 6.1, with results for VFL accuracy and communication efficiency presented in Sections 6.2 and  6.3, respectively. Additional evaluations, including scalability, training time, and performance on real datasets, are discussed in Appendix G. Each experiment elucidates results and provides relevant insights, highlighting (1) the performance-communication tradeoff of NN-based and boosting-based methods, (2) the necessity for advanced communication-efficient algorithms for imbalanced distributed datasets, and (3) the scalability potential of VFL algorithms.
@@ 6.1 Experimental Settings
This subsection includes the datasets, evaluated algorithms, and training methodology. Detailed dataset specifications, environments, and hyperparameter settings can be found in Appendix F.
Datasets. Our experimental design incorporates seven public datasets, namely covtype dat (1998), msd dat (2011), gisette dat (2008a), realsim dat (2015), epsilon dat (2008b), letter dat (1991), and radar Khosravi (2020), detailed in Appendix F. The msd dataset is used for regression tasks, while the others cater to classification tasks. Each dataset is partitioned into 80% training and 20% testing instances. The datasets’ features are distributed among multiple parties (typically four), split based on feature importance ($\alpha$) or correlation ($\beta$). In the correlation-based split, each party is assigned an equal number of dataset features.
Algorithms. We assess extensive code-available VFL algorithms in our experiments, including split-NN-based (SplitNN Vepakomma et al. (2018), C-VFL Castiglia et al. (2022a)), split-GBDT-based (FedTree Li et al. (2023), SecureBoost Cheng et al. (2021a), Pivot Wu et al. (2020a)), and ensemble-based (GAL Diao et al. (2022a)) algorithms. AL Xian et al. (2020a) is excluded due to its inferiority to GAL Diao et al. (2022a). For fairness, experiments are conducted without encryption or noise addition. In light of the reported minor variations in accuracy and communication (w/o encryption) among split-GBDT-based methods like FedTree, SecureBoost, and Pivot due to precision issues Cheng et al. (2021a); Wu et al. (2020a), we have elected to use FedTree as a representative in our evaluation of their performance and communication costs.
@@ 6.2 VFL Accuracy
In this subsection, we study the performance of VFL algorithms by varying the data split parameters, $\alpha$ and $\beta$, and assessing the resulting impact on the accuracy. Our analysis includes a range of algorithm types, namely split-NN-based, split-GBDT-based, and ensemble-based methods. The performance is detailed in Table 2. From our exploration, we can draw three key observations.
The influence of split parameters $\alpha$ and $\beta$ on VFL performance varies significantly with the choice of algorithm and dataset. The performance of certain algorithms, such as SplitNN and FedTree, remains relatively consistent across different $\alpha$ and $\beta$ values. For others, notably C-VFL, these parameter changes can cause substantial variations in performance. For instance, on the epsilon dataset, C-VFL’s accuracy fluctuates by up to 12% and 10% when $\alpha$ and $\beta$ are adjusted from 0.1 to 100 and from 0 to 1.0, respectively. Despite the potential significant influence of $\alpha$ and $\beta$ parameters, their effect on accuracy seems to be contingent upon specific dataset-algorithm combinations. This underlines the importance of extensive evaluations across a broader spectrum of $\alpha$ and $\beta$ values, a critical step towards illustrating the robustness of VFL algorithms.
@@
The compression of SplitNN-based methods, particularly when employed on imbalanced partitioned datasets, can significantly impact accuracy. While C-VFL’s model structure is akin to SplitNN, the incorporation of compression results in C-VFL having the lowest accuracy among all the tested baselines. This is particularly pronounced in cases of imbalanced importance distribution, i.e., smaller $\alpha$. For example, when $\alpha=0.1$, C-VFL’s performance on the letter and epsilon datasets is barely superior to random guessing. This highlights a pressing need for further exploration and development of compression methods suited for biased partition scenarios.
covtype SplitNN 91.2±0.4% 92.1±0.2% 92.1±0.3% 92.1±0.1% 91.7±0.3% 91.8±0.2% 92.1±0.1% 92.3±0.2%
GAL 66.1±2.9% 60.2±1.1% 62.1±4.1% 61.5±3.9% 63.9±1.2% 62.5±1.7% 61.8±1.6% 61.8±1.8%
FedTree 77.9±0.1% 77.8±0.2% 77.8±0.1% 77.8±0.1% 77.9±0.1% 77.9±0.2% 77.8±0.1% 77.8±0.1%
C-VFL 46.9±4.1% 14.8±19.0% 26.9±25.3% 38.3±18.6% 53.6±6.8% 50.1±1.8% 49.0±9.0% 49.5±1.0%
msd SplitNN 10.1±0.0% 10.2±0.0% 10.1±0.0% 10.1±0.0% 10.2±0.0% 10.1±0.0% 10.2±0.0% 10.1±0.0%
GAL 12.2±0.0% 12.2±0.0% 12.2±0.0% 12.2±0.0% 12.2±0.0% 12.2±0.0% 12.2±0.0% 12.2±0.0%
FedTree 10.4±0.0% 10.4±0.0% 10.4±0.0% 10.4±0.0% 10.4±0.0% 10.4±0.0% 10.4±0.0% 10.4±0.0%
C-VFL 12.7±0.0% 12.7±0.0% 12.7±0.0% 12.7±0.0% 12.3±0.0% 12.2±0.0% 12.1±0.1% 12.3±0.0%
gisette SplitNN 96.8±0.4% 96.8±0.3% 96.9±0.3% 96.9±0.2% 97.1±0.2% 97.2±0.4% 97.0±0.1% 96.8±0.4%
GAL 94.4±1.7% 95.7±0.6% 96.2±0.6% 96.1±0.6% 96.5±0.3% 96.9±0.2% 96.6±0.3% 96.9±0.1%
FedTree 97.0±0.3% 97.1±0.5% 97.1±0.5% 97.1±0.3% 97.0±0.4% 96.9±0.3% 97.1±0.2% 97.0±0.3%
C-VFL 91.0±11.5% 97.8±3.4% 90.1±11.8% 94.0±4.0% 90.1±8.8% 97.6±4.1% 97.6±4.1% 93.6±5.2%
realsim SplitNN 97.1±0.0% 97.0±0.0% 97.0±0.1% 97.1±0.1% 97.0±0.0% 97.0±0.2% 97.0±0.1% 97.0±0.1%
GAL 92.4±0.7% 93.5±2.2% 96.1±0.4% 96.4±0.1% 96.5±0.1% 96.5±0.0% 96.5±0.0% 96.5±0.1%
FedTree 85.1±0.2% 85.2±0.2% 85.1±0.1% 85.1±0.2% 85.2±0.2% 85.2±0.2% 85.1±0.2% 85.2±0.1%
C-VFL 89.6±17.4% 88.6±16.7% 89.5±17.4% 87.8±16.1% 89.3±17.2% 89.3±17.2% 87.8±16.2% 89.2±17.0%
epsilon SplitNN 86.3±0.1% 86.2±0.1% 86.3±0.0% 86.2±0.0% 85.9±0.1% 85.9±0.0% 86.0±0.1% 86.2±0.1%
GAL 88.2±0.6% 85.9±1.5% 85.7±0.4% 86.3±0.2% 86.1±0.0% 86.1±0.4% 86.3±0.3% 86.4±0.2%
FedTree 77.2±0.1% 77.3±0.1% 77.2±0.0% 77.2±0.1% 77.3±0.1% 77.3±0.0% 77.3±0.0% 77.3±0.1%
C-VFL 50.1±0.0% 53.5±4.9% 67.7±20.0% 61.6±16.4% 70.5±28.9% 65.8±22.3% 53.4±4.7% 60.8±15.1%
letter SplitNN 95.5±0.3% 96.0±0.3% 96.1±0.2% 96.0±0.3% 96.3±0.2% 96.0±0.3% 96.2±0.3% 96.1±0.2%
GAL 57.9±4.1% 51.4±3.1% 52.0±4.3% 49.1±2.6% 49.1±2.2% 50.0±3.0% 48.3±4.4% 48.8±2.0%
FedTree 91.7±0.3% 91.9±0.4% 92.0±0.3% 92.0±0.3% 92.0±0.2% 92.0±0.2% 91.9±0.4% 91.9±0.3%
C-VFL 5.0±1.0% 12.6±8.3% 37.1±47.1% 43.8±4.5% 34.8±45.0% 8.1±1.9% 3.6±0.7% 33.6±39.8%
GAL 90.0±3.6% 95.9±2.5% 97.5±0.4% 97.9±0.1% 96.7±0.1% 96.9±0.2% 97.6±0.4% 97.7±0.3%
FedTree 99.3±0.0% 99.3±0.0% 99.3±0.0% 99.3±0.0% 99.3±0.0% 99.3±0.0% 99.3±0.0% 99.3±0.0%
C-VFL 59.9±47.8% 53.4±38.5% 43.1±23.9% 37.3±15.7% 30.6±6.3% 53.8±39.0% 60.8±49.0% 54.4±39.9%
6.3 Communication Efficiency
Figure 3: The total communication size of VFL algorithms (50 global iterations)
In this subsection, we evaluate VFL algorithms’ communication efficiency by analyzing their total communication size within 50 fixed epochs, as shown in Figure 3. Additional communication details, such as the maximum incoming and outgoing communication, are provided in Appendix G.1. Given that FedTree, Pivot, and SecureBoost incur comparable communication costs when excluding encryption overhead, we will utilize FedTree as a representative for the other two for simplicity. Upon examining the figure, two main observations can be drawn.
Gradient-boosting algorithms, including GAL and FedTree, generally exhibit smaller communication sizes compared to neural-network-based algorithms like SplitNN and C-VFL, with the exception of the letter dataset with 26 classes. C-VFL’s compression techniques, though limiting its communication size, cannot match the efficiency of GAL and FedTree, even with a significant accuracy trade-off. The higher communication cost in neural networks is due to frequent transmission of gradients and representations, a factor that boosts SplitNN’s optimal accuracy.
Additionally, the efficiency of FedTree and GAL is contingent on the global dataset’s size. The primary distinction between FedTree and GAL lies in the type of information received on the primary party side. GAL collates prediction results from all secondary parties, with the size being proportional to the number of instances. Conversely, FedTree gathers the histogram from all secondary parties, with the size being proportional to the number of features. Consequently, GAL incurs lower communication costs than FedTree on high-dimensional datasets, such as gisette and realsim, while maintaining comparable communication costs on other datasets.

### v2

@@ 2.3 Evaluate Party Correlation
[…]
The lower bound occurs if and only if $\mathbf{X}_{1}$ comprises all features of either $\mathbf{D}_{1}$ or $\mathbf{D}_{2}$, with $\mathbf{X}_{2}$ containing the remaining features. The upper bound occurs if and only if $\mathbf{X}_{1}$ holds $m$ features from both $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, with $\mathbf{X}_{2}$ holding the remaining $m$ features from $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$.
3 Split Synthetic VFL Datasets
This section aims to develop algorithms to split features according to two key factors: importance and correlation. These algorithms should allow users to adjust the party importance and correlation of synthetic VFL datasets by simply modulating two parameters: $\mathbf{\alpha}$ and $\beta$. The intended mapping should meet two criteria: (1) The scope of $\mathbf{\alpha}$ and $\beta$ should encompass a broad spectrum of feature splits, inclusive of both real splits and random splits. (2) When two global datasets bear similarities, synthetic VFL datasets derived from them using identical $\mathbf{\alpha}$ and $\beta$ parameters should yield similar VFL algorithm behaviors. We provide both theoretical and empirical validation for criteria (1) in this section, whereas criteria (2) is substantiated through experiments in Section 4.4.
3.1 Split by Party Importance
In light of the computational expense incurred by the Shapley value method, an alternative and more efficient strategy is necessary to perform feature splits based on importance. With all parties exhibiting symmetry in the context of $\mathbf{X}$, varying the importance among parties essentially translates to varying the variance of the importance among them. Assuming each party $P_{i}$ possesses an importance factor $\alpha_{i}>0$, we propose the implementation of the Dirichlet distribution parameterized by $\mathbf{\alpha}=\{\alpha_{i}\}_{i=1}^{K}$ for feature splitting. This approach ensures two beneficial properties post-split: (1) a larger $\alpha_{i}$ guarantees a higher expected importance for $P_{i}$, and (2) a smaller $\|\{\alpha_{i}\}_{i=1}^{K}\|_{2}$ assures a greater variance in the importance among parties.
More specifically, we propose a feature splitting method based on feature importance. After initializing local datasets for each party, a series of probabilities $r_{1},\dots,r_{K}$ s.t. $\sum_{i=1}^{K}r_{i}=1$ is sampled from a Dirichlet distribution $\text{Dir}(\alpha_{1},\dots,\alpha_{K})$. Each feature is randomly allocated to a party $P_{k}$, selected based on the probabilities $r_{k}$. To accommodate algorithms that fail when faced with empty features, we can ensure each party is initially provided with a random feature before the algorithm is set in motion. Detailed formalization of this algorithm can be found in Appendix C.
Theorem 2.
[…]
@@ 3.1 Split by Party Importance
The proof of Theorem 2 can be found in Appendix A, resembling the Dirichlet-multinomial mean proof but focusing on sum importance instead of feature counts. The metric of importance, $\phi_{j}(v)$, comprises the Shapley value and the recently proposed Shapley-CMI Han et al. (2021). Theorem 2 asserts that the expected cumulative importance $\mathbb{E}[Z_{i}]$ of each party is proportional to the importance parameter $\alpha_{i}$. The Dirichlet-based split method ensures that: (1) a larger value of $\alpha_{i}$ leads to a higher expected value of $r_{i}$, thus a higher expected value of party importance, and (2) a smaller value of $\|\{\alpha_{i}\}_{i=1}^{K}\|_{2}$ results in a larger variance in $r_{i}$, as well as more imbalanced importance among parties. Both properties are empirically validated in Appendix D.2. Hence, the proposed method naturally aligns with the requirements for feature importance. With $\alpha=1$, Dirichlet-split mirrors a uniform distribution, incorporating random splits within the uniform scope. Even for manual equal splits lacking consistent criteria, a large $\alpha$ in Dirichlet-split can encapsulate them by yielding nearly equal feature distribution among parties.
3.2 Split by Party Correlation
@@
This correlation-based feature-split algorithm (Alg. 1) is designed to allocate features across multiple parties based on a given correlation parameter $\beta$. The algorithm’s operation is premised on a defined number of features for each party, represented as $m_{1},\dots,m_{K}$. Commencing with the initialization of a column permutation matrix $\mathbf{P}$ to an identity matrix (line 1), the algorithm proceeds to define a score function, $f(\mathbf{P};\mathbf{X})$, which represents the overall correlation Icor after the features are permutated by $\mathbf{P}$ (line 2). Subsequently, the algorithm determines the range of the score function (lines 3-4). This forms the basis for calculating the target correlation $f^{*}(\mathbf{X};\beta)$, which is a linear interpolation between the lower and upper bounds controlled by the correlation index $\beta$ (line 5). Next, the algorithm locates the optimal permutation matrix $\mathbf{P}^{*}$ by solving an permutation-based optimization problem. Notably, we employ the Biased Random-Key Genetic Algorithm (BRKGA) Gonçalves and Resende (2011) for this purpose. The final step of the algorithm splits the features according to the derived optimal permutation and the pre-set number of features for each party (lines 6-7).
@@
The efficiency of the optimization process, involving numerous Icor invocations, is crucial. For smaller datasets, Singular Value Decomposition (SVD) Baker (2005) is used for direct singular value computation. However, for high-dimensional datasets, we employ truncated SVD Hansen (1990) estimates the largest top-$d_{t}$ singular values, assuming the remainder as zero for standard variance calculation. The ablation study of $d_{t}$ is included in Appendix G.6. Our experiments, detailed in Appendix D.2, confirm the efficacy of both split methods.
3.3 Compare Feature Split Across Global Datasets
The metrics presented in Section 2 facilitate meaningful comparisons of feature splits within the same global datasets but fall short when comparing across different datasets. To bridge this gap and enable a comparison between real and synthetic VFL datasets, we introduce methods to map these metrics to two values: $\alpha$ and $\beta$, where $\alpha$ indicates party balance and $\beta$ indicates party correlation. Consequently, this mapping enables a direct comparison between feature splits originating from real and synthetic VFL datasets, as demonstrated in Figure 0(b).
To estimate $\alpha$, the importance of each party is calculated by Shapley values. These importance are then normalized and treated as Dirichlet parameters $\alpha_{i}$ for each party $P_{i}$, in line with Theorem 2. To approximate the scale of the Dirichlet parameters and align them with the generation of synthetic datasets, we find a symmetric Dirichlet distribution $\text{Dir}({\alpha})$ that has the same variance as $\text{Dir}(\alpha_{1},\dots,\alpha_{K})$, as given in Proposition 4. This value of ${\alpha}$ reflects the variance of party importance. The proof is provided in Appendix A.
@@ 4 Experiment
This section benchmarks cutting-edge VFL algorithms, with a detailed review provided in Section 4.1. Experimental settings are outlined in Section 4.2, and results regarding VFL accuracy and synthetic-real correlation are in Sections 4.3 and 4.4, respectively. Further evaluations, such as real communication cost, scalability, training time, and real dataset performance, are in Appendix G. Each experiment elucidates results and provides relevant insights, highlighting (1) the performance-communication tradeoff of NN-based and boosting-based methods, (2) the performance similarity between synthetic and real VFL datasets under the same $\alpha,\beta$, and (3) the scalability potential of VFL algorithms.
4.1 Review of VFL Algorithms
This section critically reviews current VFL algorithms, with a focus on accuracy, efficiency, and communication size. VertiBench concentrates on standard supervised learning tasks such as classification and regression within synchronized parties, summarized in Table 1. Notably, this benchmark excludes studies exploring other aspects Jin et al. (2021); Qi et al. (2022); Jiang et al. (2022) and other tasks Chang et al. (2020); Li et al. (2021b); Chen and Zhang (2022); He et al. (2022); Li et al. (2022b). While most VFL algorithms presume accurate inter-party data linking, we adopt this approach in VertiBench, despite recent contrary findings Wu et al. (2022a); Nock et al. (2021) that this assumption may not be true. We refer to parties with and without labels as primary and secondary parties respectively.
Table 1: Summary of existing VFL algorithms
Category Model1 Algorithm Contribution Reference Data2 Feature3
Ensemble- based Any AL Accuracy Xian et al. (2020) Syn Manual
GAL Accuracy Diao et al. (2022) Syn Manual
Split- based NN SplitNN Accuracy Vepakomma et al. (2018) Syn N/A
C-VFL Communication Castiglia et al. (2022) Syn Manual
BlindFL Efficiency Fu et al. (2022b) Syn Manual
FedOnce Communication Wu et al. (2022b) Syn Random
GBDT SecureBoost Accuracy Cheng et al. (2021) Syn Manual
Pivot Accuracy Wu et al. (2020) Syn Manual
FedTree Accuracy, Efficiency Li et al. (2023) Syn Random
VF2Boost Efficiency Fu et al. (2021) Syn Manual
RF Fed-Forest Communication Liu et al. (2020) Syn Random
  • 1 Abbreviations: NN - neural network; GBDT - gradient boosting decision trees; RF - random forest; Any - model-agnostic.
  • 2 Dataset in experiments: Syn - synthetic datasets partitioned from global datasets.
  • 3 Datasets used in the experiments: Manual - features manually split without specific reasons; Random - features randomly split without explanation; N/A - no VFL experiments conducted.
Most of the existing VFL methods can be categorized into ensemble-based and split-based. Ensemble-based methods have each party maintain a full model for local prediction and use collaborative ensemble techniques during training. Conversely, split-based methods delegate each party with a portion of the model, representing different inference stages. A comprehensive comparison is in Appendix B. In this paper, we concentrate on the primary types of VFL, acknowledging that there are various subtypes as identified in Liu et al. (2022). Exploring these subtypes in depth will be an objective of our future research efforts.
In our experiments, we evaluate various VFL algorithms, including split-NN-based (e.g., SplitNN, C-VFL, FedOnce), split-GBDT-based (FedTree), and ensemble-based (GAL). For fairness, evaluations exclude encryption or noise. Noting minor variances among split-GBDT-based methods such as FedTree and SecureBoost, FedTree is used as a representative in our experiments.
@@ 4.2 Experimental Settings
This subsection includes the datasets and training method. Detailed dataset specifications, environments, and hyperparameter settings can be found in Appendix F.
Datasets. Our experiments utilize 11 datasets: nine centralized ones (covtype Blackard (1998), msd Bertin-Mahieux (2011), gisette Guyon et al. (2008), realsim Andrew (2015), epsilon Guo-Xun et al. (2008), letter Slate (1991), radar Khosravi (2020), MNIST Deng (2012), CIFAR10 Krizhevsky and Hinton (2009)), and two real-world VFL datasets (NUS-WIDE Chua et al. (2009), Vehicle Duarte and Hu (2004)), with detailed descriptions available in Appendix F. The msd dataset is used for regression tasks, while the others cater to classification tasks. Each dataset is partitioned into 80% training and 20% testing instances except NUS-WIDE, MNIST, and CIFAR10 with pre-defined test set. The datasets’ features are distributed among multiple parties (typically four), split based on party importance ($\alpha$) or correlation ($\beta$). In the correlation-based split, each party is assigned an equal number of features.
@@ 4.3 VFL Accuracy
In this subsection, we assess the impact on the performance of VFL algorithms when varying $\alpha$ and $\beta$. Our analysis includes all the three VFL categories in Table 1. The performance is summarized in Figure 3 and detailed in Table 9 in Appendix G. The result on msd dataset provides similar insights to others, thus only included in Table 9. From our exploration, we can draw three key observations.
Split parameters $\alpha$ and $\beta$ significantly affect VFL algorithm performance, depending on the algorithm and dataset. SplitNN and FedTree show stable performance across various $\alpha$ and $\beta$ settings. In contrast, C-VFL demonstrates notable performance fluctuations: up to 10% on epsilon and 40% on letter with varying $\alpha$. GAL performs better on imbalanced datasets (affected by $\alpha$ by 8% on letter and radar, 2-5% on others) and is minimally influenced by $\beta$. FedOnce, favoring balanced and highly correlated datasets, is affected by $\alpha$ (5-10% on letter, gisette, epsilon) and by $\beta$ (1-3% on covtype, epsilon). These findings highlight the need for comprehensive evaluations across a range of $\alpha$ and $\beta$ to determine VFL algorithms’ robustness.
@@
The compression of SplitNN renders them particularly affected by party imbalance. C-VFL, modelled after SplitNN, exhibits the least accuracy among tested baselines due to its compression approach. Moreover, C-VFL exhibits marked sensitivity to the imbalance level, $\alpha$. Specifically, at $\alpha=0.1$, its performance on datasets like letter and epsilon scarcely surpasses random guessing. However, C-VFL thrives in highly imbalanced split of radar dataset. This data-dependent behavior underscores an urgent need to refine compression techniques for VFL tailored to varying imbalances.
Figure 3: Accuracy of VFL algorithms on different datasets varying imbalance and correlation
4.4 Performance Correlation: VertiBench Scope vs. Real Scope
In assessing the performance correlation between VertiBench-synthetic and real VFL datasets, we use derived $\alpha$ and $\beta$ values of NUS-WIDE and Vehicle (Section 3.3) to generate comparable synthetic datasets. To evaluate the relative performance of each algorithm, we calculate the accuracy differences between Vehicle-synthetic and NUS-WIDE-synthetic datasets for each algorithm and compare with real dataset accuracy differences, with further details in Appendix G.8.
Our experiment reveals a positive correlation between relative algorithm performance on synthetic datasets with matching $\alpha$ and $\beta$, and their performance on real VFL datasets. This indicates that, under the same $\alpha$ or $\beta$, higher mean accuracy on synthetic datasets typically implies better performance on real VFL datasets, thus affirming the relevance of VertiBench-synthetic datasets in approximating real VFL performance.
Figure 4: Mean accuracy differences: synthetic datasets vs. real datasets